Opened 5 years ago

Closed 4 years ago

#523 closed flight processing (fixed)

GB13/10, flight day 156/2014, Wytham Woods

Reported by: knpa Owned by: benj
Priority: alpha 4 medium Milestone: 2014 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by tec)

Data location: /data/visuyan2/scratch/arsf/2014/flight_data/uk/GB13_10-2014_156_Wytham/

Data arrived from ARSF via network transfer on 05/06/2014

GB13/10 Scientific objective: unknown

GB13/10 Priority: 8

PI: Anne Verhoef

Notes: There is OWL data for this flight. There is FW LiDAR for this flight.

Sensors:

Camera (Requested but lower priority), delivered 21/08/2014, Archvied 13/05/2015
Fenix (Requested), delivered 21/08/2014, Archvied 13/05/2015
LiDAR (Requested), delivered 21/08/2014, Archvied 13/05/2015
FW LiDAR, delivered 21/08/2014
OWL

Change History (31)

comment:1 Changed 5 years ago by knpa

  • Description modified (diff)

comment:2 Changed 5 years ago by stgo

Tipo didn't leave any notes, as such I am collecting what I can from the processing cfgs. He has however created a delivery for the hyperspectral data. I will submit this for delivery check once a readme is prepared.
SCT values below:

-1 0.95
-2 -0.1
-3 0.95
-4 0.94
-5 0.95
-6 0.94
-7 0.94
-8 0.94
-9 0.95
-10 43.95
-11 35.95
-12 0.95
-13 0.94

Pixel size used: 1.2

Inspection of the data presents a couple of cloud shadows on the data but otherwise it seems fine.

comment:3 Changed 5 years ago by stgo

Created new readme, ready for D/C

comment:4 Changed 5 years ago by tec

Started RCD Processing

comment:5 Changed 5 years ago by tec

RCD images not in the LiDAR boundary:
26, 68, 82, 96, 124, 174, 192

Last edited 5 years ago by tec (previous) (diff)

comment:6 Changed 5 years ago by tec

Finished RCD processing. Awaiting delivery check.

comment:7 Changed 5 years ago by dap

Completed delivery checking processed RCD data. Awaiting second delivery check from stgo.

comment:8 Changed 5 years ago by dap

Some dark photographs in the delivery due to clouds, a comment should be made about this in the read me file.

comment:9 Changed 5 years ago by tec

Starting delivery check.

comment:10 Changed 5 years ago by tec

Line 12 header file is missing.
flightlines/mapped/f156123b_mapped_osng.bil.hdr

No mosaic with vectors, check with PI if they want vectors.

apl_corr command is wrong, thinks it wants to use NEXTMAP dem files not ASTER dem files.

apl_corr is also wrong in the readme, its using absolute paths so contains filepaths with visuyan in it, it also had NEXTMAP dems in the command.

Readme says UNKNOWN SENSOR not fenix just after apl_corr command.

fenix_viewvector_list.txt doesnt have windows line endings, check if they need to be the windows style of CRLF

In the doc directory there are the CSS files
ISO019115-stylesheet.xsl
ISO019115-css.css

Level 1 line 1 bands 356-368 look dark and has overflows. Band 621 is blank.

Have not checked the other level 1 lines or the level 3 mapped files.

comment:11 Changed 5 years ago by stgo

Currently processing the LiDAR data. Roll error of -0.005552596 seems to be consistent across flightlines. Now classifying.

comment:12 Changed 5 years ago by stgo

Lidar with full waveform created ready for DC.

comment:13 Changed 5 years ago by stgo

Made corrections noted above and added readme comment about overflows. REady for re D/C

*also note I've emailed the PI about waiting on vectors.

Last edited 5 years ago by stgo (previous) (diff)

comment:14 Changed 5 years ago by tec

Added in the readme about underflows. Ready to deliver. Underflow bands 490-540 590-621. Running zip_mapped.sh.

comment:15 Changed 5 years ago by dap

Beginning LiDAR delivery check

comment:16 Changed 5 years ago by dap

LiDAR delivery check complete, the following issues should be addressed before delivery of the data:

  • No full waveform files found
  • If las1.2 files were created, will need to change read me as it says 1.0 currently.
  • LiDAR ASCII files don't have CRLF terminators - should they?
  • There is a .tif file in the screenshots directory - should this be there?
  • One of the lines appears to be of an odd shape (LDR-GB13_10-2014-156-07.LAS), which wasn't mentioned in the data quality remarks.

comment:17 Changed 5 years ago by stgo

Lidar ready for recheck.

comment:18 Changed 5 years ago by dap

Rechecking LiDAR delivery.

comment:19 Changed 5 years ago by dap

LiDAR Delivery Check

The following issues were found and need to be resolved:

  • FW LAS files have been given names that don't adhere to the file naming convention.
  • GRIMM data was acquired, but not highlighted in the logsheet.
  • Data quality remarks uses word "ditched" - this is slang.
  • enp files have been left in the dem directory, these should be removed.

The following issues were found and have been resolved:

  • full-waveform.pdf missing from doc directory (copied from day 169's LiDAR delivery directory).
  • Intensity tif has been left in screenshots (removed)

PI is happy to wait for vectors, but would like to be kept informed of progress. Missing vectors screenshot is therefore not a problem.

comment:20 Changed 5 years ago by dap

Checked LiDAR delivery issues indicated above after corrections were made. No issues found. LiDAR now ready for delivery.

comment:21 Changed 5 years ago by dap

Creating delivery for all data from day 156.

comment:22 Changed 5 years ago by dap

All data dispatched to Prof. Verhoef, 21/08/2014.

comment:23 Changed 5 years ago by dap

  • Component changed from Processing: general to Archiving
  • Description modified (diff)
  • Owner set to benj

comment:24 Changed 5 years ago by tec

Fenix
Changed line 2 SCT to -0.08

comment:25 Changed 5 years ago by tec

Fenix
Ready for delivery check.

comment:26 Changed 5 years ago by knpa

Done Fenix delivery check.

All looks good apart from:

Minor problems:

  • Fodis dir should not be present

Don't forget to zip up the mapped files.

comment:27 Changed 5 years ago by tec

Fenix
Removed fodis directory.

Zipping mapped files now.

comment:28 Changed 5 years ago by tec

Fenix
Delivery available on FTP arsf4 for PI

comment:29 Changed 4 years ago by tec

Archiving
Starting archiving

comment:30 Changed 4 years ago by tec

  • Description modified (diff)

Archiving
Done confirmed as of 13/05/2015 12:10

comment:31 Changed 4 years ago by dac

  • Resolution set to fixed
  • Status changed from new to closed
Note: See TracTickets for help on using tickets.